New Georgia-Pacific lumber plant in Albany, GA is expected to produce 300 million board-feet of lumber a year. The facility will employ 125 employees in the production section working in 8-hour shifts each day. Assume number of work-days in the facility is 300. a. What is the productivity per-employee? 
 b. What is the productivity per labor-hour? 
Answer #1

Output = 300 million = 300000000 per year

There are 125 employees working per day and there are 300 days in a year. So number of employees per year = 125 × 300 = 37500 employees

Each employee works 8 hour per day. So for 37500 employees total labor hour per year = 37500 × 8 = 300000 hours

a) Productivity per employee = output/number of employees = 300000000/37500 = 8000 per employee

b) productivity per labor hour = output/labor hour per year = 300000000/300000 = 1000 per labor hour